OpenStudy (anonymous):

The rectangle below has a total perimeter of 125 cm.

OpenStudy (anonymous):

Write an equation and solve it to find the length of the longer side of the rectangle. (1 point) 2x + 10 = 125; x = 57.5 cm x + 5 = 125; x = 120 cm 5x = 125; x = 25 cm 25x = 125; x = 5 cm

OpenStudy (smored):



OpenStudy (anonymous):

that's not one of the options though how do I make it one?

OpenStudy (anonymous):

It would be option A right?!

OpenStudy (anonymous):

@Nurali

OpenStudy (anonymous):

@CGGURUMANJUNATH

OpenStudy (smored):

2x+2(5)=125 is 2x+10=125